OpenStudy (anonymous):

can you help me please I need help with these 4 questions two times one and four−fifths one−half divided by six two divided by three−eighths two−thirds times three and three−eighths

OpenStudy (anonymous):

I only need help with two divided by three−eighths

OpenStudy (anonymous):

well 3/8 equals= 0.375, so try dividing 2 by that number and see what you get

OpenStudy (anonymous):

hmmm I got a really big number for some reason.. I got 5.333

OpenStudy (anonymous):

I need it in a faction

OpenStudy (anonymous):

in a fraction? oh, hold on just a sec then letme figure this out. Doyou have any options to choose from, or fill in?

OpenStudy (anonymous):

its five and one−third or two and one−fourth

OpenStudy (anonymous):

okay hold on just one sec.. I think I know what to do jus tmaking sure I"m doing it right

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

Yes, I'm still trying to figure this out.. I don't hink i've ever done this before

OpenStudy (phi):

two divided by three−eighths \[ 2 \div \frac{3}{8} \] to divide a number (or a fraction) by a fraction, flip the 2nd fraction and multiply \[ 2 \div \frac{3}{8} \text{ becomes } 2 \cdot \frac{8}{3} \] you get \(\frac{16}{3}\) or \( 5\frac{1}{3}\) or \(5.\bar3\) If you think of a ruler, you might see there are a lot of 3/8 ths of an inch in 2 inches

OpenStudy (anonymous):

okay thanks!!

OpenStudy (anonymous):

okay